# Open dynamical systems with a moving hole

### v1

@@ 6. Final remarks
(ii) The survivor set $K^{\omega}$ studied in this paper involves one moving hole. Our results can be extended to study survivor sets with multiple moving holes for sufficiently large $b$ and $m$. More precisely, for $m,M\in\mathbb{N}$, let $\vec{\omega}=(\omega(1),\omega(2),\ldots,\omega(M))\in\prod_{\ell=1}^{M}(D_{b}^{m})^{\mathbb{N}}$, where each $\omega(\ell)=\omega^{0}(\ell)\omega^{1}(\ell)\ldots\in\left(D_{b}^{m}\right)^{\mathbb{N}}$. Then the survivor set with multiple moving holes $\vec{\omega}$ is defined by
@@
:=$\mathcal{F}:=\left\{f_{a}(x)=rO(x+a):a\in\mathcal{D}\right\},$
where $O$ is an orthogonal matrix in $\mathbb{R}^{d}$ so that $\mathcal{F}$ satisfies the strong separation condition. For $\omega=\omega^{0}\omega^{1}\ldots\in(\mathcal{D}^{m})^{\mathbb{N}}$, let
@@ Acknowledgements
The authors thank Natalia Jurga for bringing their attention to the finiteness property conjecture of joint spectral radius of real matrices. The authors would also thank Dejun Feng for some useful discussions on the multifractal results in Remark 1.7 (i). Moreover, the authors thank Aihua Fan for his suggestion on Remark 1.2. The first author was supported by Chongqing NSF: CQYC20220511052 and Scientific Research Innovation Capacity Support Project for Young Faculty No. ZYGXQNISKYCXNLZCXM-P2P. The third author was supported by NSFC No. 12471085 and the China Postdoctoral Science Foundation No. 2024M763857.

### v2

@@ 7. Final remarks
(ii) The survivor set $K^{\omega}$ studied in this paper involves one moving hole. Our results can be extended to study survivor sets with multiple moving holes for sufficiently large $b$ and $m$. More precisely, for $m,M\in\mathbb{N}$, let $\vec{\omega}=(\omega(1),\omega(2),\ldots,\omega(M))\in\prod_{\ell=1}^{M}(D_{b}^{m})^{\mathbb{N}_{0}}$, where each $\omega(\ell)=\omega^{0}(\ell)\omega^{1}(\ell)\ldots\in\left(D_{b}^{m}\right)^{\mathbb{N}_{0}}$. Then the survivor set with multiple moving holes $\vec{\omega}$ is defined by
@@
$\mathcal{F}:=\left\{f_{a}(x)=rO(x+a):a\in\mathcal{D}\right\},$
where $O$ is an orthogonal matrix in $\mathbb{R}^{d}$ so that $\mathcal{F}$ satisfies the strong separation condition. For $\omega=\omega^{0}\omega^{1}\ldots\in(\mathcal{D}^{m})^{\mathbb{N}_{0}}$, let
@@ Acknowledgements
The authors thank the anonymous referee for many useful comments which greatly improve the paper, in particular Theorem 1.8. The authors would also thank Natalia Jurga for bringing their attention to the finiteness conjecture of real matrices, and thank Hanfeng Li and Dejun Feng for some useful discussions. Moreover, the authors thank Aihua Fan for his suggestion on Remark 1.2. The first author was supported by Chongqing NSF: CQYC20220511052 and Scientific Research Innovation Capacity Support Project for Young Faculty No. ZYGXQNISKYCXNLZCXM-P2P. The third author was supported by NSFC Nos. 12501110, 12471085 and the China Postdoctoral Science Foundation No. 2024M763857.